OpenStudy (anonymous):

Verify that y is an explicit solution to the differential equation, y''+y=tan(x) , y =-cos(x)ln(secx+tanx).... just help with differentiation

OpenStudy (anonymous):

that's cheating dawg

OpenStudy (anonymous):

wouldn't it be product rule with chain

OpenStudy (anonymous):

this is what my teach will put on our test lol for sure... feel my pain

OpenStudy (experimentx):

yeah ... you can do that ... use product rule to differentiate ... then again differentiate again ... it will take lot of time!!

OpenStudy (anonymous):

yes but we don't have wolfram alpha at test time haha anyways i'll do that for a test problem next week

OpenStudy (experimentx):

yeah .. best of luck!!

jimthompson5910 (jim_thompson5910):

y =-cos(x)ln(secx+tanx) y' = d/dx[ -cos(x)ln( sec(x) + tan(x) ) ] y' = d/dx[ -cos(x)]*ln( sec(x) + tan(x) ) -cos(x)*d/dx[ln( sec(x) + tan(x) )] y' = sin(x)*ln( sec(x) + tan(x) ) - cos(x)*(1/( sec(x) + tan(x) ))*d/dx[ sec(x) + tan(x) ] y' = sin(x)*ln( sec(x) + tan(x) ) - cos(x)*(1/( sec(x) + tan(x) ))*(sec(x)tan(x) + sec^2(x)) y' = sin(x)*ln( sec(x) + tan(x) ) - cos(x)*(1/( sec(x) + tan(x) ))*(sec(x)(tan(x) + sec(x))) y' = sin(x)*ln( sec(x) + tan(x) ) - cos(x)*sec(x) y' = sin(x)*ln( sec(x) + tan(x) ) - 1 Now repeat this to find y'' y'' = d/dx[ sin(x)*ln( sec(x) + tan(x) ) - 1 ] y'' = d/dx[ sin(x)*ln( sec(x) + tan(x) )] - d/dx[ 1 ] y'' = d/dx[ sin(x)*ln( sec(x) + tan(x) )] y'' = d/dx[ sin(x)]*ln( sec(x) + tan(x) ) + sin(x)*d/dx[ln( sec(x) + tan(x) ) ] y'' = cos(x)*ln( sec(x) + tan(x) ) + sin(x)*(1/(sec(x) + tan(x)) )*d/dx[sec(x) + tan(x)] y'' = cos(x)*ln( sec(x) + tan(x) ) + sin(x)*(1/(sec(x) + tan(x)) )*(sec(x)*tan(x) + sec^2(x)) y'' = cos(x)*ln( sec(x) + tan(x) ) + sin(x)*(1/(sec(x) + tan(x)) )*(sec(x)(tan(x) + sec(x))) y'' = cos(x)*ln( sec(x) + tan(x) ) + sin(x)*sec(x) y'' = cos(x)*ln( sec(x) + tan(x) ) + tan(x) =========================================== y''+y=tan(x) ( cos(x)*ln( sec(x) + tan(x) ) + tan(x) )+ ( -cos(x)ln(secx+tanx) )=tan(x) tan(x) = tan(x)

jimthompson5910 (jim_thompson5910):

It's a lot of work, but once you get the hang of derivatives and you've mastered them (and perhaps memorized them), then it shouldn't be too bad.